# adF: a Novel System for Measuring Web Fingerprinting through Ads

### v2

@@ Abstract.
This paper introduces adF, a novel system for analyzing the vulnerability of different devices, Operating Systems (OSes), and browsers to web fingerprinting. adF performs its measurements from code inserted in ads. We have used our system in several ad campaigns that delivered 5.40 million ad impressions. The collected data allow us to assess the vulnerability of current desktop and mobile devices to web fingerprinting. Based on our results, we estimate that 66% of desktop devices and 40% of mobile devices can be uniquely fingerprinted with our web fingerprinting system. However, the resilience to web fingerprinting varies significantly across browsers and device types, with Chrome on desktops being the most vulnerable configuration.
@@
To counter web fingerprinting, we propose ShieldF, a simple solution which blocks the reporting by browsers of those attributes that we found in the analysis of our dataset that present the most significant discrimination power. Our experiments reveal that ShieldF outperforms all anti-fingerprinting solutions proposed by major browsers (Chrome, Safari and Firefox) offering an increase in the resilience offered to web fingerprinting up to 62% for some device configurations. ShieldF is available as an add-on for any chromium-based browser. Moreover, it is readily adoptable by browser and mobile app developers. Its widespread use would lead to a significant improvement in the protection offered by browsers and mobile apps to web fingerprinting.
fingerprinting, advertising, user privacy, adTag, web tracking, browser developers
copyright: nonedoi: isbn: copyright: noneconference: TBD; ;
1. Introduction
@@
Digital advertising is arguably the primary source of funding for the current Internet. Some of the most important tech companies, such as Google or Meta (previously Facebook), declare that over 77,5% and 97,5% (2023 data) of their profits come from digital advertising, respectively (advertising_revenue_2023_google, ; advertising_revenue_2023_meta, ).
@@
The appeal of online advertising for advertisers, compared to traditional forms of publicity (e.g., TV, radio, or newspapers), lies in its ability to deliver personalized ads. Implementing personalized advertising is based on techniques that identify unique users online. These techniques are used for tracking, profiling users, displaying targeted or re-targeted ads, implementing attribution models, etc. The most widely used techniques for this purpose are third-party cookies, advertising IDs, and fingerprinting (cahn2016empirical, ; al2018beyond, ; bujlow2017survey, ).
@@
Third-party cookies and advertising IDs, when properly implemented with consent acquisition (not always the case), comply with data protection regulations. Fingerprinting, by definition, is an intrusive practice and represents a severe threat to end-users’ privacy. The most common type of fingerprinting involves embedding a script in a website. When a user visits such a website, the script collects attributes from the browser, OS, and device rendering the web page. The concatenation of these attributes creates a device fingerprint. If the number of attributes is large enough, the fingerprint can make the device unique even among hundreds of thousands or millions of devices. This practice is known as web fingerprinting. The privacy risks associated with web fingerprinting have prompted extensive research in this area (mayer2009any, ; eckersley2010, ; mowery2012pixel, ; nikiforakis2013cookieless, ; acar2013fpdetective, ; acar2014web, ; laperdrix2016beauty, ; englehardt2016online, ; kaizer2016towards, ; gomezBoix2018hiding, ; cao2017cross, ; olejnik2017battery, ).
A body of work proposes countermeasure techniques to prevent fingerprinting (laperdrix2015mitigating, ; Al_FannahNasserMohammed2020Tltl, ; azad2020taming, ; canvasbloquer_firefox, ; wu2019rendered, ; laperdrix2016beauty, ; iqbal2021fingerprinting, ; privacybadger, ; faizkhademi2015fpguard, ; nikiforakis2015privaricator, ; laperdrix2017fprandom, ; trickel2019everyone, ; gomez2019collaborative, ; obidatcanvas2020, ; salomatin2022comparison, ; fernandez2023keep, ; torok2023only, ; noscript, ). Browser developers (e.g., Apple, Mozilla, or Google) have also proposed anti-fingerprinting measures in later releases of their browsers (cameron_2018_appleagainstfp, ; enhanced_tracking_protection_mozillawiki, ; chrome_plans_fp_ua, ). However, a rapid web technology development has made some existing proposals obsolete and requires a continuous effort from browser developers to update their anti-fingerprinting solutions.
@@
In this rapidly evolving context, proper auditing methodologies are crucial for measuring the risk of web fingerprinting to users and assessing vulnerabilities in different web browsers. The research community has proposed measurement solutions in this direction (eckersley2010, ; laperdrix2016beauty, ; boda2011user, ; cao2017cross, ; browserleaks, ; privacycheck, ; fietkau2021fpmon, ). Essentially, these solutions involve deploying a script on a web page to compute the device fingerprint when a device connects to the page. This methodology assesses whether the computed fingerprint is unique within the database of previously computed fingerprints. This approach has been very useful in improving our understanding of web fingerprinting. However, it presents some limitations. On the one hand, if the auditing script is deployed on a website meant for measuring fingerprinting (eckersley2010, ; eckersley_panopticlick, ; laperdrix2016beauty, ; laperdrix_amiunique, ), it requires the voluntary participation of users who are willing to assess their fingerprint. Sustaining a continuous flow of visitors to these types of website over time poses a challenge. On the other hand, running the script on a regular website (gomezBoix2018hiding, ; andriamilanto2021largescale, ; li2020wtmbf, ) introduces biases in the type of devices/users visiting that single website. Moreover, the ability to conduct such experiments depends on the willingness of the website owner to participate. Furthermore, the coverage of this technique is limited to web-browsers, leaving aside other very common venues such as mobile apps that operate on top of web views, essentially utilizing web technologies.
@@
In this paper, we propose a novel measurement technology to study web fingerprinting. Our methodology utilizes ads, which are HTML elements included on web-pages and within mobile apps. We embed a script to compute the device fingerprint within these ads, overcoming limitations of existing solutions: 1) We run ad campaigns that perform measurements transparently to users, ensuring longitudinal analyses without relying on volunteers; 2) Our methodology allows for the study of web fingerprinting within the context of mobile apps; 3) Our data sample is not limited to measurements from a single vantage point (i.e., web page); 4) Researchers have full control over our technique, deciding when to run the campaigns; 5) By harnessing the targeting capabilities of online advertising, our methodology enables the definition of specific device configurations (based on device type, OS, and browser) to be measured for the first time.
@@
We implement our methodology in a system we refer to as adF (as discussed in Section 4) and conducted real ad campaigns that collected more than 4,04M fingerprint samples in three independent periods of time over the course of a year and a half (as detailed in Section 5). First, these real experiments demonstrate that our methodology achieves an accuracy exceeding 80%, confirming its validity for the study of web fingerprinting. Second, the analysis of the collected fingerprint samples allows to assess the vulnerability of different device configurations to web fingerprinting. We define a device configuration as the combination of device type (desktop vs. mobile), an operating system (e.g., Windows, macOS, Android, or iOS), and a browser (e.g., Mozilla Firefox, Chrome, or Safari). In the case of mobile apps, we only distinguish between operating systems, Android vs. iOS (see Section 8).
Our analysis reveals that 64% of desktop devices can be uniquely fingerprinted from ads rendered by browsers. In the case of mobile devices, 40% (35%) can be fingerprinted with ads delivered to browsers (mobile apps). Moreover, we observe varying degrees of vulnerability to fingerprinting among different device configurations. For instance, in mobile devices, iOS offers significantly better protection against web fingerprinting compared to Android. In the case of desktop devices, over 78% of devices running Chrome can be fingerprinted, regardless of the operating system. This percentage drops to around 5,5% for {Safari, macOS} and stands at 22% for {Firefox, Linux}.
@@
Finally, a thorough analysis of all attributes contributing to web fingerprinting reveals that cardinality (number of distinct values an attribute can take) and entropy (frequency of appearance of the different values) are the two crucial parameters that define the discrimination power of attributes. The device configurations most vulnerable to fingerprinting are those associated with a higher number of attributes with significant discrimination power (as discussed in Section 9). Building on this insight, we propose ShieldF a straightforward approach, easily adoptable by browser developers, to counter web fingerprinting. ShieldF is a browser extension compatible with all Chromium-based browsers. It prevents the reporting of attributes with high discrimination power to third-parties while preserving the user’s browsing experience. Our experiments show that ShieldF outperforms most popular countermeasure defenses in the literature, and the anti-fingerprinting solutions of all major browser developers (Section 10), enhancing the resilience to web fingerprinting across all considered device configurations by up to 62%.
2. Background
2.1. Identifying individual users
@@
Advertising companies use different techniques to identify unique users and target them based on individual preferences such as user demographics, location, and interests. Although they can generally only identify devices rather than users (as in the open advertising ecosystem of websites and mobile apps, there is no registration process), they assume that a single user uses a device. According to (cahn2016empirical, ; al2018beyond, ; bujlow2017survey, ), the most common techniques used to identify unique users are: third-party cookies, advertising IDs, and web fingerprinting techniques.
2.1.1. Third-party cookies
@@ 2.1.2. Advertising ID
The advertising ID (Ad-ID) is a unique user identifier from the UUID standard (uuid_wiki, ) assigned to an electronic device (smartphone, tablet, PC) and is intended for advertising purposes. It allows advertisers to track users’ advertising activity to personalize their offers (advertising_ID_wiki, ) through browsers and mobile applications. The Ad-ID is embedded in the ad requests received by ad-tech stakeholders, which use the Ad-ID to enrich the information known about such Ad-ID, e.g., demographic information or interests associated to the Ad-ID. The Ad-ID is available in all major desktop and mobile operating systems (Windows, iOS, macOS, Android, etc.) and usually receives a different name for each operating system, for instance, AAID (Android Advertising ID) in Android or IDFA (Identifier for Advertisers) in Apple’s operating systems.
2.1.3. Web fingerprinting
@@
Web browsers have developed a set of APIs that allow accessing different parameters of the browser, the operating system, or the device. The original goal of these functions is to optimize the functionality of browsers and the user experience. For instance, one of these functions exposes the device’s screen size such that the web page’s content can be appropriate for the specific screen size.
@@
Web fingerprinting is implemented through a script embedded in the HTML code of the web page. This script leverages the different functions and APIs offered by the browser to collect different attributes from the browser configuration (e.g., browser add-ons or configured languages), the OS (installed fonts), and the device (sound device or graphic card information). In conjunction, the values of all these attributes create a fingerprint of the device. If a sufficiently large number of attributes are collected, the resulting fingerprint of the device might be unique even among a large pool of devices.
@@
Note that, contrary to the case of third-party cookies or advertising IDs, which require implicit or explicit consent from the user, fingerprinting is implemented without any type of consent from the user (fp_guidance_w3c, ) and thus it is the most privacy-aggressive practice of the three. Nonetheless, web fingerprinting also has legit uses, such as pursuing fraud or credential theft, by verifying the legitimacy of users when logging into specific sites (moalosi2019combating, ; gabryel2020browser, ; liu2022gummy, ; durey2021fp, ; laperdrix2019morellian, ; durey2022theses, ).
2.2. Web measurements from ads
@@
The online advertising industry has evolved into an intricate ecosystem for measuring user engagement, drawing on data from JavaScript scripts embedded within ads delivered to websites and mobile apps. These scripts, belonging to different third-party companies, exploit browser APIs to collect data, which is then processed and sent to third-parties’ back-end servers. These scripts can be used to measure KPIs (e.g., the click-through ratio of an ad campaign), detect fraudulent practices, and more. In this paper, we use this same measurement technology to implement our script to study web fingerprinting. We embed a script in an ad (rather than a web page) and gather information about the browser configuration, OS, and device software and hardware while aiming to identify a user uniquely. Nonetheless, note scripts executed within ads face limitations compared to those running natively in a website’s HTML code due to security measures like the Cross-Domain Policy (crossdomainpolicy_acrobatguide, ; websecurity_sameoriginpolicy_mdn, ).
2.2.1. Web fingerprinting from ads
@@
One of the main uses of web fingerprinting is to reconstruct the web browsing history of a user. A tracking entities embeds their fingerprinting script in as many websites as possible to gather data about the user’s browsing history. It can also be used for re-targeting, i.e., users are shown ads about products they previously viewed on other websites or e-commerce sites. However, web fingerprinting may not be effective in these tasks since it is not feasible to show the same ad to a user on every page they visit. Nonetheless, web fingerprinting from ads can be very useful in building attribution models for advertisers. An attribution model is an extensively used and very valuable tool used by advertisers to understand the performance of their ad campaigns, web-pages, and ad channels. To build this model, advertisers must reconstruct each user’s path. This is a vector including all the ads shown to an individual user and the interaction of the user with them. By analyzing the paths of thousands of users, a model can be constructed to infer the most successful elements (ad creatives, ad campaigns, web-pages, mobile apps, etc) that lead to more conversions, such as purchases.
@@
To create users’ paths, it is crucial to uniquely identify each user. In this context, web fingerprinting from ads is a plausible alternative to building attribution models. Even with the imminent discontinuation of third-party cookies (blocked_chrome_3partycookie_2024, ; blocked_apple_3partycookie_2020, ), and the significant restrictions imposed on the use of the advertising ID (advertisingID_announcementGoogle, ; idfa_changes_ios14, ) which limits the number of devices disclosing it [^1] Just 48,52% of the 4.04M samples collected in the dataset used in this paper reported the Ad-ID., it may become the only practical approach to building attribution models in the near future.
@@
2.3. Popular browser’ approaches to counter fingerprinting

### v3

@@ Abstract
This paper introduces adF, a novel system for analyzing the vulnerability of different devices, operating systems, and browsers to web fingerprinting. adF performs its measurements from code inserted in ads. We use our system in several ad campaigns that delivered 5.40 million ad impressions. The collected data allow us to assess the vulnerability of current desktop and mobile devices to web fingerprinting. We estimate that 66% of desktop and 40% of mobile devices can be uniquely fingerprinted with our web fingerprinting system. However, web fingerprinting resilience varies significantly across browsers and device types, with Chrome on desktops being the most vulnerable configuration.
@@
To counter web fingerprinting, we propose ShieldF, a straightforward solution that blocks browsers from reporting attributes with the most significant discrimination power, as identified through our dataset analysis. Our experiments reveal that ShieldF outperforms all anti-fingerprinting solutions proposed by major browsers (Chrome, Safari and Firefox), providing a resilience boost of up to 62% for some device configurations. ShieldF is accessible as an add-on for any Chromium-based browser. Moreover, it is readily adoptable by browser and mobile app developers. Its widespread use would lead to a marked enhancement in browsers’ and mobile apps’ defense against web fingerprinting.
I Introduction
@@
Digital advertising is arguably the primary source of funding for the current Internet. Some of the most important tech companies, such as Google or Meta (previously Facebook), declare that over 77.5% and 97.5% (2023 data) of their profits come from digital advertising [advertising_revenue_2023_google, advertising_revenue_2023_meta], respectively.
@@
The appeal of online advertising for advertisers, compared to traditional forms of publicity (e.g., TV, radio, or newspapers), lies in its ability to deliver personalized ads. Implementing personalized advertising is based on techniques that identify unique users online. These techniques are used for tracking, profiling users, displaying targeted or re-targeted ads, implementing attribution models, etc. The most widely used techniques for this purpose are third-party cookies, advertising IDs, and fingerprinting [bujlow2017survey].
@@
Third-party cookies and advertising IDs, when properly implemented with consent acquisition (not always the case), comply with data protection regulations. Fingerprinting, by definition, is an intrusive practice and represents a severe threat to end-users’ privacy. The most common type of fingerprinting involves embedding a script in a website. When a user visits such a website, the script collects attributes from the browser, OS, and device rendering the web-page. The concatenation of these attributes creates a device fingerprint. If the number of attributes is large enough, the fingerprint can make the device unique even among hundreds of thousands or millions of devices. This practice is known as web fingerprinting. The privacy risks associated with web fingerprinting have prompted extensive research in this area [mayer2009any, eckersley2010, acar2014web, laperdrix2016beauty, gomezBoix2018hiding, cao2017cross]. A body of work proposes countermeasure techniques to prevent fingerprinting [laperdrix2015mitigating, Al_FannahNasserMohammed2020Tltl, azad2020taming, canvasbloquer_firefox, wu2019rendered, laperdrix2016beauty, iqbal2021fingerprinting, privacybadger, nikiforakis2015privaricator, laperdrix2017fprandom, noscript]. Browser developers (e.g., Apple, Mozilla, or Google) have also proposed anti-fingerprinting measures in later releases of their browsers [cameron_2018_appleagainstfp, enhanced_tracking_protection_mozillawiki, user_agent_reduction_privacysandbox]. However, a rapid web technology development has made some existing proposals obsolete and requires a continuous effort from browser developers to update their anti-fingerprinting solutions.
@@
In this rapidly evolving context, proper auditing methodologies are crucial for measuring the risk of web fingerprinting to users and assessing vulnerabilities in different web browsers. The research community has proposed measurement solutions in this direction [eckersley2010, laperdrix2016beauty, boda2011user, cao2017cross]. Essentially, these solutions involve deploying a script on a web-page to compute the device fingerprint when a device connects to the page. This methodology assesses whether the computed fingerprint is unique within the database of previously computed fingerprints. This approach has been very useful in improving our understanding of web fingerprinting. However, it presents some limitations. On the one hand, if the auditing script is deployed on a website meant for measuring fingerprinting [eckersley2010, laperdrix2016beauty], it requires the voluntary participation of users who are willing to assess their fingerprint. Sustaining a continuous flow of visitors to these types of website over time poses a challenge. On the other hand, running the script on a regular website [gomezBoix2018hiding, andriamilanto2021largescale, li2020wtmbf] introduces biases in the type of devices/users visiting that single website. Moreover, the ability to conduct such experiments depends on the willingness of the website owner to participate. Furthermore, the coverage of this technique is limited to web-browsers, leaving aside other very common venues such as mobile apps that operate on top of WebViews, essentially utilizing web technologies.
@@
In this paper, we propose a novel measurement technology to study web fingerprinting. Our methodology utilizes ads, which are HTML elements included on web-pages and within mobile apps. We embed a script to compute the device fingerprint within these ads, overcoming limitations of existing solutions: 1) We run ad campaigns that perform measurements transparently to users, ensuring longitudinal analyses without relying on volunteers; 2) Our methodology allows for the study of web fingerprinting within the context of mobile apps; 3) Our data sample is not limited to measurements from a single vantage point (i.e., web-page); 4) Researchers have full control over our technique, deciding when to run the campaigns; 5) By harnessing the targeting capabilities of online advertising, our methodology enables the definition of specific device configurations (based on device type, OS, and browser) to be measured for the first time.
@@
We implement our methodology in a system we refer to as adF (as discussed in Section IV) and conducted real ad campaigns that collected more than 4.04M fingerprint samples in three independent periods of time over the course of a year and a half (as detailed in Section V). First, these real experiments demonstrate that our methodology achieves an accuracy exceeding 80%, confirming its validity for the study of web fingerprinting. Second, the analysis of the collected fingerprint samples allows to assess the vulnerability of different device configurations to web fingerprinting. We define a device configuration as the combination of device type (desktop vs. mobile), an operating system (e.g., Windows, macOS, Android, or iOS), and a browser (e.g., Mozilla Firefox, Chrome, or Safari). In the case of mobile apps, we only distinguish between operating systems, Android vs. iOS (see Section VII).
@@
Our analysis reveals that 66% of desktop devices can be uniquely fingerprinted from ads rendered by browsers. In the case of mobile devices, 40% (35%) can be fingerprinted with ads delivered to browsers (mobile apps). Moreover, we observe varying degrees of vulnerability to fingerprinting among different device configurations. For instance, in mobile devices, iOS offers significantly better protection against web fingerprinting compared to Android. In the case of desktop devices, over 78% of devices running Chrome can be fingerprinted, regardless of the operating system. This percentage drops to around 5.5% for {Safari, macOS} and stands at 22% for {Firefox, Linux}. Finally, a thorough analysis of all attributes contributing to web fingerprinting reveals that cardinality (number of distinct values an attribute can take) and entropy (frequency of appearance of the different values) are the two crucial parameters that define the discrimination power of attributes. The device configurations most vulnerable to fingerprinting are those associated with a higher number of attributes with significant discrimination power (as discussed in Section IX). Building on this insight, we propose ShieldF a straightforward approach, easily adoptable by browser developers, to counter web fingerprinting. ShieldF is a browser extension compatible with all Chromium-based browsers. It prevents the reporting of attributes with high discrimination power to third-parties while preserving the user’s browsing experience. Our experiments show that ShieldF outperforms most popular countermeasure defenses in the literature, and the anti-fingerprinting solutions of all major browser developers (Section X), enhancing the resilience to web fingerprinting across all considered device configurations by up to 62%.
II Background
II-A Identifying individual users
@@
Advertising companies use different techniques to identify unique users and target them based on individual preferences such as user demographics, location, and interests. Although they can generally only identify devices rather than users (as in the open advertising ecosystem of websites and mobile apps, there is no registration process), they assume that a single user uses a device. According to [bujlow2017survey], the most common techniques used to identify unique users are: third-party cookies, advertising IDs, and web fingerprinting techniques.
II-A1 Third-party cookies
@@ II-A2 Advertising ID
The advertising ID (Ad-ID) is a unique user identifier from the UUID standard assigned to an electronic device (smartphone, tablet, PC) and is intended for advertising purposes. It allows advertisers to track users’ advertising activity to personalize offers through browsers and mobile apps. The Ad-ID is embedded in the ad requests received by ad-tech stakeholders, which use the Ad-ID to enrich the information known about such Ad-ID, e.g., demographic information or interests associated to the Ad-ID. The Ad-ID is available in all major desktop and mobile operating systems (Windows, iOS, macOS, Android, etc.) and usually receives a different name for each operating system, for instance, AAID (Android Advertising ID) in Android or IDFA (Identifier for Advertisers) in Apple’s operating systems.
II-A3 Web fingerprinting
@@
Web browsers have developed a set of APIs that allow accessing different parameters of the browser, the operating system, or the device. The original goal of these functions is to optimize the functionality of browsers and the user experience. For instance, one of these functions exposes the device’s screen size such that the web-page’s content can be appropriate for the specific screen size.
@@
Web fingerprinting is implemented through a script embedded in the HTML code of the web-page. This script leverages the different functions and APIs offered by the browser to collect different attributes from the browser configuration (e.g., browser add-ons or configured languages), the OS (installed fonts), and the device (sound device or graphic card information). In conjunction, the values of all these attributes create a fingerprint of the device. If a sufficiently large number of attributes are collected, the resulting fingerprint of the device might be unique even among a large pool of devices.
@@
Note that, contrary to the case of third-party cookies or Ad-IDs, which require implicit or explicit consent from the user, fingerprinting is implemented without any type of consent from the user and thus it is the most privacy-aggressive practice of the three. Nonetheless, web fingerprinting also has legit uses, such as pursuing fraud or credential theft, by verifying the legitimacy of users when logging into specific sites [gabryel2020browser].
II-B Web measurements from ads
@@
The online advertising industry has evolved into an intricate ecosystem for measuring user engagement, drawing on data from JavaScript scripts embedded within ads delivered to websites and mobile apps. These scripts, belonging to different third-party companies, exploit browser APIs to collect data, which is then processed and sent to third-parties’ back-end servers. These scripts can be used to measure KPIs (e.g., the click-through ratio of an ad campaign), detect fraudulent practices, and more. In this paper, we use this same measurement technology to implement our script to study web fingerprinting. We embed a script in an ad (rather than a web-page) and gather information about the browser configuration, OS, and device software and hardware while aiming to identify a user uniquely. Nonetheless, note scripts executed within ads face limitations compared to those running natively in a website’s HTML code due to security measures like the Cross-Domain Policy.
II-B1 Web fingerprinting from ads
@@
One of the main uses of web fingerprinting is to reconstruct the web browsing history of a user. Tracking entities embed their fingerprinting script in as many websites as possible to gather data about the user’s browsing history. It can also be used for re-targeting, i.e., users are shown ads about products they previously viewed on other websites or e-commerce sites. However, web fingerprinting may not be effective in these tasks since it is not feasible to show the same ad to a user on every page they visit. Nonetheless, web fingerprinting from ads can be very useful in building attribution models for advertisers. An attribution model is an extensively used and very valuable tool used by advertisers to understand the performance of their ad campaigns, web-pages, and ad channels. To build this model, advertisers must reconstruct each user’s path. This is a vector including all the ads shown to an individual user and the interaction of the user with them. By analyzing the paths of thousands of users, a model can be constructed to infer the most successful elements (ad creatives, ad campaigns, web-pages, mobile apps, etc) that lead to more conversions, such as purchases.
@@
To create users’ paths, it is crucial to uniquely identify each user. In this context, web fingerprinting from ads is a plausible alternative to building attribution models. Even with the imminent discontinuation of third-party cookies [blocked_chrome_3partycookie_2024], and the significant restrictions imposed on the use of the Ad-ID [advertisingID_announcementGoogle] which limits the number of devices disclosing it [^1] Just 48.52% of the 4.04M samples collected in the dataset used in this paper reported the Ad-ID., it may become the only practical approach to building attribution models in the near future.
@@
II-C Popular browser’ approaches to counter fingerprinting